OpenStudy (anonymous):

Write an equation for the line that is parallel to the given line and that passes through the given point. y = 3/4 x – 9; (–8, –18) @jim_thompson5910

OpenStudy (anonymous):

\[y=\frac{ 3 }{ 4 }x-9 (-8,-18)\]

jimthompson5910 (jim_thompson5910):

What is the slope of y = 3/4 x – 9

OpenStudy (anonymous):

m=3/4 ?

jimthompson5910 (jim_thompson5910):

So the slope of any line parallel to this one will also be 3/4

jimthompson5910 (jim_thompson5910):

If two lines are parallel, then they have equal slopes

jimthompson5910 (jim_thompson5910):

\[\Large y = mx + b\] \[\Large y = \frac{3}{4}x + b\] \[\Large -18 = \frac{3}{4}(-8) + b\] solve for b

OpenStudy (anonymous):

b=-12

jimthompson5910 (jim_thompson5910):

good, so your equation is \[\Large y = \frac{3}{4}x - 12\]